# Ar2Can: An Architect and an Artist Leveraging a Canvas for Multi-Human Generation

### v1

@@ 4.2 Results
Simple
Matching
Hungarian
Centroid
Matching
Curriculum
Training
Count
Multi-ID
HPS
@@ B.2 Models
Architect-A uses Qwen-2.5-0.5B. Architect-B uses Flux-Schnell (1B parameters, 4 steps). The Artist is a Flux-Kontext (12B parameters). We use RetinaFace [deng2020retinaface] (ResNet-50 backbone) for face detection with confidence threshold 0.5. To compute ArcFace embeddings [deng2019arcface], we use the Anetlopev2111https://github.com/deepinsight/insightface model, producing 512-dimensional features. For pose computation, we use the rtmlib library  [^2] https://github.com/Tau-J/rtmlib with their default model.
@@ B.4 Inference
20:     PuLID [guo2024pulid] (pose/expression variation), or
@@
UMO-UNO [cheng2025umo]: We obtained the code from the official repository [^4] https://github.com/bytedance/UMO and used the default configuration with 25 inference steps and default guidance scale as recommended by the authors.
X-Verse [chen2025xverse]: We obtained the code from the official repository [^5] https://github.com/bytedance/XVerse and followed the default settings with 28 inference steps and default guidance scale.
UMO-OmniGen2 [cheng2025umo]: We obtained the code from the official repository [^6] https://github.com/VectorSpaceLab/OmniGen and used the recommended configuration with 50 inference steps and default guidance scale.
Dream-O [mou2025dreamo]: We obtained the code from the official repository [^7] https://github.com/bytedance/DreamO and adopted the recommended Flux.1-turbo variant with 12 inference steps and default guidance scale as specified in the documentation.
ID-Patch [zhang2025id]: We obtained the code from the official repository [^8] https://github.com/bytedance/ID-Patch and used all default parameters from the official implementation without modification.
WithAnyone [xu2025withanyone]: We obtained the code from the official repository [^9] https://github.com/Doby-Xu/WithAnyone and used the default parameters except for the SigLIP weight, which we set to 0.5 as we found it to be the optimal trade-off between identity preservation and prompt alignment.
@@ C.1 Canvas Construction Pipeline
We start the process by generating a DisCo [disco2025] scene with $N$ people and detecting face locations (Stage 1-2). For generating the canvas in a pose-controlled training, we additionally estimate human pose skeletons from the DisCo image. Then, we initialize a blank canvas and replace each synthetic face with a reference face from our curated sources (Stage 3-4). In pose mode, we overlay the estimated pose skeletons onto the canvas, providing additional spatial guidance for action and pose alignment.
Figure C.1: Illustration of the Canvas Construction Pipeline. Left: Image generated by DisCo [disco2025], using the input prompt. Middle: Standard canvas. Each canvas combines reference faces sampled from different data sources: $\mathcal{D}_{1}$ (multi-view references with secondary views), $\mathcal{D}_{2}$ (single-view references with PuLID-generated or augmented secondary views), and $\mathcal{D}_{3}$ (the original synthetic face from DisCo). Right: If we want to perform pose-mode training, we overlay the human pose skeletons estimated from the DisCo images. We make sure to blend with lesser intensity in face bbox regions.
@@ C.2 Frontal Pose Reward Formulation
We design a lightweight 2D frontal pose reward using 5-point facial landmarks (left eye, right eye, nose, left mouth, right mouth) detected by RetinaFace [deng2020retinaface]. For each detected face $i$, we compute a frontality score $\delta_{i}$ combining roll angle and yaw asymmetry:
@@ D.3 Grid Search on Reward Weights
Count
$\alpha$
HPSv3
$\beta$
Face
$\zeta$
Pose
$\eta$
Count Acc. Multi-ID Prompt Align

### v2

@@ 4.2 Results
Simple Matching Hungarian Centroid Matching Curriculum Training Count Multi-ID HPS
@@ B.2 Models
Architect-A uses Qwen-2.5-0.5B. Architect-B uses Flux-Schnell (1B parameters, 4 steps). The Artist is a Flux-Kontext (12B parameters). We use RetinaFace [9] (ResNet-50 backbone) for face detection with confidence threshold 0.5. To compute ArcFace embeddings [10], we use the Anetlopev2111https://github.com/deepinsight/insightface model, producing 512-dimensional features. For pose computation, we use the rtmlib library  [^2] https://github.com/Tau-J/rtmlib with their default model.
@@ B.4 Inference
20:     PuLID [15] (pose/expression variation), or
@@
UMO-UNO [8]: We obtained the code from the official repository [^4] https://github.com/bytedance/UMO and used the default configuration with 25 inference steps and default guidance scale as recommended by the authors.
X-Verse [7]: We obtained the code from the official repository [^5] https://github.com/bytedance/XVerse and followed the default settings with 28 inference steps and default guidance scale.
UMO-OmniGen2 [8]: We obtained the code from the official repository [^6] https://github.com/VectorSpaceLab/OmniGen and used the recommended configuration with 50 inference steps and default guidance scale.
Dream-O [23]: We obtained the code from the official repository [^7] https://github.com/bytedance/DreamO and adopted the recommended Flux.1-turbo variant with 12 inference steps and default guidance scale as specified in the documentation.
ID-Patch [42]: We obtained the code from the official repository [^8] https://github.com/bytedance/ID-Patch and used all default parameters from the official implementation without modification.
WithAnyone [38]: We obtained the code from the official repository [^9] https://github.com/Doby-Xu/WithAnyone and used the default parameters except for the SigLIP weight, which we set to 0.5 as we found it to be the optimal trade-off between identity preservation and prompt alignment.
@@ C.1 Canvas Construction Pipeline
We start the process by generating a DisCo [5] scene with $N$ people and detecting face locations (Stage 1-2). For generating the canvas in a pose-controlled training, we additionally estimate human pose skeletons from the DisCo image. Then, we initialize a blank canvas and replace each synthetic face with a reference face from our curated sources (Stage 3-4). In pose mode, we overlay the estimated pose skeletons onto the canvas, providing additional spatial guidance for action and pose alignment.
Figure C.1: Illustration of the Canvas Construction Pipeline. Left: Image generated by DisCo [5], using the input prompt. Middle: Standard canvas. Each canvas combines reference faces sampled from different data sources: $\mathcal{D}_{1}$ (multi-view references with secondary views), $\mathcal{D}_{2}$ (single-view references with PuLID-generated or augmented secondary views), and $\mathcal{D}_{3}$ (the original synthetic face from DisCo). Right: If we want to perform pose-mode training, we overlay the human pose skeletons estimated from the DisCo images. We make sure to blend with lesser intensity in face bbox regions.
@@ C.2 Frontal Pose Reward Formulation
We design a lightweight 2D frontal pose reward using 5-point facial landmarks (left eye, right eye, nose, left mouth, right mouth) detected by RetinaFace [9]. For each detected face $i$, we compute a frontality score $\delta_{i}$ combining roll angle and yaw asymmetry:
@@ D.3 Grid Search on Reward Weights
Count $\alpha$ HPSv3 $\beta$ Face $\zeta$ Pose $\eta$ Count Acc. Multi-ID Prompt Align